OpenStudy (anonymous):

x^2+y^2=25 36x^2-4y^2=144

OpenStudy (anonymous):

x^2+y^2=25 9x^2-y^2=36 + -------------- 10x^2=61

OpenStudy (anonymous):

x =\[\pm \sqrt{6.1}\]

OpenStudy (anonymous):

the same for y :)